OpenStudy (fibonaccichick666):

In need of a quick hw check: Find an ex. of a discontinuous fn s.t. \(f:[0,1]\rightarrow R\) where the Intermediate Value Theorem fails.

OpenStudy (fibonaccichick666):

So I defined \[f(x)=[0 ~if~ x\in Q ~and~ 0\le x\le 1, ~~1~if~x\not{\in} Q ~and~0\le x\le 1]\] then for any y s.t. 0<y<1 the IVT fails. Does this work? Is it actually answering the question?

OpenStudy (fibonaccichick666):

@jim_thompson5910 , have you had analysis?

ganeshie8 (ganeshie8):

the first function i think of is 1/(1-2x)

OpenStudy (fibonaccichick666):

why 2x?

ganeshie8 (ganeshie8):

this function has no zeroes eventhough it changes sign in (0,1)

OpenStudy (fibonaccichick666):

i was just thinking my answer doesn't fit the problem and was at 1/(.5-x)

ganeshie8 (ganeshie8):

i think we want discontinuity somewhere between (0,1) right

OpenStudy (fibonaccichick666):

well, it could be discontinuous anywhere

OpenStudy (fibonaccichick666):

but the issue is thinking of a fn that maps 0,1 onto all of the reals

ganeshie8 (ganeshie8):

Oh you want an onto function is it ?

OpenStudy (fibonaccichick666):

yea, my function doesn't satisfy the conditions I don't think

OpenStudy (zzr0ck3r):

your function is fine for the conditions stated, but it is NOT surjective.

OpenStudy (fibonaccichick666):

well, I mean the conditions make it have to be surjective, don't they?

OpenStudy (zzr0ck3r):

why? \(f:A\rightarrow B\) does not imply \(f\) is onto \(B\) it just implies \(B\) is the codomain.

OpenStudy (fibonaccichick666):

You do not read that as [0,1] is mapped onto the reals?

OpenStudy (zzr0ck3r):

nope

OpenStudy (fibonaccichick666):

oh, that's how I read it

OpenStudy (zzr0ck3r):

http://en.wikipedia.org/wiki/Function_(mathematics) scroll down to notation. Sometimes people use a double arrow to indicate onto, but there is no standard, because we use that same double arrow to mean "the trivial map".

OpenStudy (fibonaccichick666):

...hmm I have always read it as onto. Maybe not spanning now that I think of it though

OpenStudy (fibonaccichick666):

alright, so I can use this. Thanks for the clarification. I have another which as a technicality is bothering me, can I tag you?

OpenStudy (zzr0ck3r):

Unless your book/teacher uses that notation to be onto, but I have never seen that. Notice also that @ganeshie8 also did not assume surjectivity.

OpenStudy (fibonaccichick666):

true ok I guess I was just so used to reading it as onto for some reason

OpenStudy (zzr0ck3r):

it would make more sense... more bad notation in mathematics.

OpenStudy (fibonaccichick666):

true, well hopefully it works because I just can't think of a disc. fn with that domain which could span the reals...unless it was like a point disc